# ELDeR: Getting Efficient LLMs through Data-Driven Regularized Layer-wise Pruning

### v1

@@ 4.2 Performance Comparison
For fairness, we randomly selected 128 samples from the WikiText-2 training dataset with a sequence length of 2048 as the data for ELDeR to get layer weights and perform regularization, as well as the calibration data for the baselines. Following [15], we selected another 1,000 samples from WikiText-2 training dataset to perform RFT leveraging LoRA [58] on the baselines after pruning. The aforementioned 128 samples and the 1,000 samples have no overlap. Because ELDeR is finetuning-free, we did not perform RFT on ELDeR. The pruning ratio was 25% which means 25% of the transformer layers in a model will be pruned.
@@
Table 1: Performance comparison of different pruning methods. ‘PPL’ refers to perplexity on WikiText-2. The accuracy is reported on five zero-shot benchmarks. The best result is in bold, and the second-best is underlined. ELDeR-$\ell_{2}$ denotes the use of the $\ell_{2}$-norm, ELDeR-$\ell_{1}$ denotes the use of the $\ell_{1}$-norm. ‘PR’ is the pruning ratio.
@@
Model Method PR PPL ($\downarrow$) PIQA(%) WinoGrande(%) HellaSwag(%) ARC-e(%) ARC-c(%) Avg_Acc(%) Phi-2 Dense 0% 5.28 79.11 75.77 73.83 78.32 54.18 72.24 SLEB 25% 7.65 68.85 63.63 50.96 49.38 30.79 52.72 ShortGPT 25% 7.15 69.67 65.19 51.26 52.46 33.89 54.49 LaCo 25% 7.38 68.53 63.76 50.39 51.28 33.45 53.48 Shortened LLaMA 25% 7.74 66.88 62.19 51.45 51.47 32.66 52.93 ELDeR-$\ell_{2}$ 25% 6.53 71.35 67.62 55.84 52.26 35.75 56.56 ELDeR-$\ell_{1}$ 25% 6.58 71.42 67.13 56.05 52.18 35.79 56.51 OPT-2.7B Dense 0% 12.46 74.81 61.01 60.58 54.42 31.14 56.39 SLEB 25% 15.71 65.19 56.26 44.54 46.28 25.14 47.48 ShortGPT 25% 14.96 67.37 57.65 46.82 49.43 26.54 49.56 LaCo 25% 15.38 66.54 59.45 43.68 48.74 24.96 48.67 Shortened LLaMA 25% 15.89 63.14 57.36 43.57 47.62 25.88 47.51 ELDeR-$\ell_{2}$ 25% 13.18 70.54 60.27 46.35 51.59 27.36 51.22 ELDeR-$\ell_{1}$ 25% 13.12 70.65 60.13 46.24 51.76 27.55 51.27 LLaMA2-7B Dense 0% 5.47 79.11 69.06 75.99 74.58 46.25 69.00 SLEB 25% 9.63 65.22 63.38 55.51 56.39 33.46 54.79 ShortGPT 25% 8.89 66.75 66.26 57.14 58.93 36.42 57.10 LaCo 25% 9.14 69.45 65.31 52.67 55.73 34.89 55.61 Shortened LLaMA 25% 9.47 65.58 64.72 58.36 54.19 32.96 55.16 ELDeR-$\ell_{2}$ 25% 7.08 72.48 67.52 60.45 62.69 39.73 60.57 ELDeR-$\ell_{1}$ 25% 7.17 72.08 67.93 60.42 62.38 39.89 60.54 LLaMA3-8B Dense 0% 5.76 85.56 77.94 79.27 78.84 56.49 75.62 SLEB 25% 10.38 72.74 64.12 67.74 65.84 45.16 63.12 ShortGPT 25% 9.26 75.38 69.25 70.12 68.54 47.57 66.17 LaCo 25% 10.14 74.69 67.52 66.36 69.43 46.31 64.86 Shortened LLaMA 25% 9.84 73.62 68.73 68.55 66.16 43.39 64.09 ELDeR-$\ell_{2}$ 25% 7.84 77.25 71.63 72.26 71.49 49.58 68.44 ELDeR-$\ell_{1}$ 25% 7.68 77.36 71.33 72.82 70.75 49.66 68.38 OPT-13B Dense 0% 10.12 76.82 64.80 69.81 61.87 35.67 61.79 SLEB 25% 11.96 72.83 64.06 63.32 59.98 34.65 58.97 ShortGPT 25% 11.38 73.59 64.52 65.68 60.41 34.99 59.84 LaCo 25% 11.79 73.96 63.24 62.17 61.46 33.28 58.82 Shortened LLaMA 25% 11.62 71.26 63.57 66.29 58.47 33.83 58.68 ELDeR-$\ell_{2}$ 25% 10.45 74.15 64.47 68.82 61.55 35.23 60.84 ELDeR-$\ell_{1}$ 25% 10.32 74.58 64.37 68.45 61.29 34.87 60.71 LLaMA2-13B Dense 0% 4.88 80.47 72.22 79.39 77.48 49.23 71.76 SLEB 25% 7.08 68.31 66.86 57.12 62.19 38.45 58.59 ShortGPT 25% 6.79 73.26 68.37 62.25 67.54 43.38 62.96 LaCo 25% 7.14 72.35 65.78 59.43 65.36 42.73 61.13 Shortened LLaMA 25% 6.92 69.45 66.38 59.86 65.25 39.12 60.01 ELDeR-$\ell_{2}$ 25% 5.82 74.56 69.34 64.79 71.25 45.63 65.11 ELDeR-$\ell_{1}$ 25% 5.89 75.06 69.18 64.96 71.43 45.26 65.18
@@
As shown in Table 1, ELDeR achieves the lowest perplexity and the highest average accuracy across all models compared to the baseline methods, demonstrating its superior performance on both generation and zero-shot tasks. On LLaMA2-7B, its perplexity is 20% lower than the second-best method, ShortGPT. Notably, on OPT-13B, compared to the dense model, ELDeR only drop 1% in average accuracy. This further demonstrates the effectiveness of ELDeR. The performance difference between $\ell_{2}$-norm and $\ell_{1}$-norm is tiny, indicating that ELDeR is insensitive to the choice of regularization norm. In the following sections, we refer to ELDeR as the version that uses the $\ell_{2}$-norm.
@@ 4.3 Acceleration
Language processing in LLMs consists of two main stages: prompt processing, which is compute-bound, and token generation, which is memory-bound. We analyze the speedup achieved in each stage separately. Table 2 shows the throughput and latency results for OPT-13B and LLaMA2-13B, evaluated on a single 80GB NVIDIA A100 GPU. Following the methodology of previous work [16], the token generation test involves generating 128-token sentences with a batch size of 64, while prompt processing latency is measured using an input sequence of 2048 tokens.
Table 2: Throughput (tokens/s) and latency (ms) on OPT-13B and LLaMA2-13B. ‘PPL’ refers to the perplexity on Wikitext2. ‘TI’ is Throughput Increase.
Model Method Pruning ratio PPL($\downarrow$) Avg_Acc(%) Tokens/s($\uparrow$) TI($\uparrow$) Latency($\downarrow$) Speedup($\uparrow$) OPT-13B Dense 0% 10.12 61.79 1029 1.00× 386.5 1.00× ELDeR 25% 10.45 60.84 1348 1.31× 286.3 1.35× ELDeR 50% 15.38 50.83 1801 1.75× 208.9 1.85× LLaMA2-13B Dense 0% 4.88 71.76 1066 1.00× 396.9 1.00× ELDeR 25% 5.82 65.11 1386 1.30× 298.4 1.33× ELDeR 50% 11.28 57.57 1823 1.71× 216.9 1.83×
At a 50% pruning ratio on OPT-13B, ELDeR achieves a 75% increase in throughput and a 46% reduction in latency compared to the dense model. For LLaMA2-13B, it delivers a 71% improvement in throughput and a 45% decrease in latency. These results underscore ELDeR’s superior efficiency in accelerating model inference.
@@ 4.4 Robustness to Different Pruning Ratios
Keeping all other settings consistent with Section 4.2, we extend the pruning ratio from 20% to 60%. The perplexity of LLaMA2-7B under different methods are shown in Figure 4. ELDeR significantly outperforms other methods across various pruning ratios. At a pruning ratio of 20%, ELDeR performs similarly to the dense model. When the pruning ratio is up to 60%, SLEB collapses, while ELDeR maintains relatively low perplexity compared to other pruning methods. This demonstrates that ELDeR maintains robust performance even under extreme pruning ratios, enabling structured pruning to unlock significant potential for model acceleration. More detailed results are in Appendix E.
@@
4.5 Dependency on Datasets
Since ELDeR relies on data-driven regularization, we investigate its dataset dependency. Keeping all other settings consistent with Section 4.2, we only change the source of the 128 samples: they are drawn respectively from Alpaca, WikiText-2, PTB, and C4. We evaluated perplexity of five methods on WikiText-2. As shown in Figure 4, ELDeR consistently outperforms the other methods across datasets, demonstrating its robustness.
4.6 Low Overhead
We evaluated the perplexity of baselines on LLaMA2-7B using varying amounts of data for RFT, keeping all other conditions consistent with Section 4.2 and only change the data size from 1,000 to 8,000. ELDeR is finetuing-free. From Figure 5, it can be observed that ELDeR outperforms the other methods with approximately 4.000 RFT data, significantly reducing data overhead. We speculate the reason is that ELDeR iteratively learns layer weights and then applies regularization, allowing it to more accurately identify layers to prune before pruning, and subsequently transfer the information from those layers to the remaining parts of the model through regularization. This process reduces information loss caused by pruning, thus preserving model performance and lowering the cost associated with RFT.
@@ 4.7 Ablation Study
Effect of Learning Layer Weights Iteratively ELDeR learns layer weights in a greedy and iterative manner. As shown in Table 3 (Row 3 and Row 6), replacing iterative layer weight learning with a one-shot approach leads to increased model perplexity and decreased accuracy, highlighting the importance of learning layer weights iteratively, which will be discussed in detail in Section 4.8.
@@
Keeping all other settings consistent with Section 4.2, we evaluate the model’s performance by varying $\lambda_{1}\in[10^{-5},10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}]$ and $\lambda_{2}\in[10^{-5},10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}]$. We perform a grid search over $\lambda_{1}$ and $\lambda_{2}$, with details provided in Appendix C. The optimal combination yielding the lowest perplexity on LLaMA2-7B is $\lambda_{1}=5\times 10^{-3}$ and $\lambda_{2}=10^{-3}$.
Figure 5: The comparison of perplexity on WikiText-2 between ELDeR and other baseline methods under different amounts of RFT data.
Table 3: Ablation results on LLaMA2-7B and LLaMA2-13B. ‘w/o W’ refers learning layer weights in one-shot, ‘w/o R’ refers no regularization. The pruning ratio is 25%.
Model Setting PPL($\downarrow$) $\Delta$ AVG_ACC $\Delta$($\downarrow$) LLaMA2-7B ELDeR 7.08 0.00 60.57 0.00 w/o W 9.26 +2.18 56.19 -4.38 w/o R 10.15 +3.07 54.36 -6.21 LLaMA2-13B ELDeR 5.82 0.00 65.11 0.00 w/o W 8.35 +2.53 59.36 -5.75 w/o R 9.47 +3.65 56.25 -8.86
@@ 4.8 Choice of Learning Layer Weights
In this section, we explore the effectiveness of (1) iteratively learning layer weights using a greedy strategy compared to (2) acquiring all layer weights at one-shot. We set the pruning ratio to 25% and conduct experiments on LLaMA2-7B (32 layers), keeping all settings consistent with Section 4.2 except for the method used to determine layer weights.
@@
4.9 Impact of Regularization
We keep other settings consistent with Section 4.2. After learning the layer weights iteratively, we consider two scenarios: (1) without regularization and (2) with regularization, followed by pruning. The perplexity and average accuracy under (1) and (2) on LLaMA2-7B are shown in Table 5. The model exhibits lower perplexity and higher average accuracy with regularization. Since perplexity is the exponential form of cross-entropy loss, lower perplexity corresponds to a lower cross-entropy loss. A smaller loss function indicates better model performance indicating that the regularization process mitigates the impact of pruning on the overall model.
According to previous work [59], the output and input of each layer in LLMs exhibit high similarity. This inherent similarity enables us to apply regularization on the difference between the input and output of certain layers using only a small portion of data. We compute the cosine similarity according to Equation 6 on the LLaMA2-7B model between the input representations $\mathbf{X_{\text{in}}^{i}}\in\mathbb{R}^{b\times n\times d}$ and the output representations $\mathbf{X_{\text{out}}^{i}}\in\mathbb{R}^{b\times n\times d}$, where $b$ denotes the batch size, $n$ the number of tokens, and $d$ is the hidden size of the model. The computation is performed on layers [3, 6, 9, 11, 13, 20, 23, 29], which are selected using the iterative method described in Section 4.8, both before and after regularization.
$\text{CosSim}(\mathbf{X_{\text{in}}^{i}},\mathbf{X_{\text{out}}^{i}})=\frac{1}{bn}\sum_{k=1}^{b}\sum_{j=1}^{n}\frac{\left\langle\mathbf{X_{\text{in}}^{i}}[k,j,:],\ \mathbf{X_{\text{out}}^{i}}[k,j,:]\right\rangle}{\left\|\mathbf{X_{\text{in}}^{i}}[k,j,:]\right\|_{2}\cdot\left\|\mathbf{X_{\text{out}}^{i}}[k,j,:]\right\|_{2}}$ (6)
As shown in Figure 7, the input-output similarity of these layers is already high before regularization and increases even further after regularization. The increased similarity in the layers with regularization indicates that the input undergoes less change after passing through these layers, suggesting that less information is retained in them. In the Appendix F, we illustrate the changes in similarity between the input and output of the layers that were not regularized, before and after regularization. The results in Figure 8 of Appendix F show that the similarity in these layers decreases, meaning that, after regularization, the input undergoes greater changes when passing through these layers than before. This demonstrates that the regularization process weakens the influence of the regularized layers and enhances the influence of the unregularized parts, suggesting that regularization process may facilitate the transfer of information from the regularized layers to the rest of the model.
Table 5: The performance differences between (1) and (2).

### v2

@@ 4.2 Performance Comparison
To ensure fairness, 128 sequences of length 2048 were randomly drawn from the WikiText-2 training set for TRSP’s two-stage regularization and the calibration process of the baselines. Following Ashkboos et al. (2024), we selected another 1,000 samples from WikiText-2 training dataset to retrain leveraging LoRA Hu et al. (2022) on the baselines after pruning. The aforementioned 128 samples and the 1,000 samples have no overlap. Because TRSP is retraining-free, we did not retrain TRSP. The pruning ratio was 25% which means 25% of the transformer layers in a model will be removed.
@@
Table 1: Performance comparison of TRSP and baselines. ‘PR’ is the pruning ratio. ‘PPL’ is the perplexity on WikiText-2. The accuracy is evaluated on five zero-shot benchmarks. TRSP -$\ell_{2}$ means using the $\ell_{2}$-norm in the second stage regularization, TRSP -$\ell_{1}$ is using the $\ell_{1}$-norm. The best result is in bold, the second-best is underlined.
@@
Model Method PR PPL ($\downarrow$) PIQA(%) WinoGrande(%) HellaSwag(%) ARC-e(%) ARC-c(%) Avg_Acc(%) Phi-2 Dense 0% 5.28 79.11 75.77 73.83 78.32 54.18 72.24 SLEB 25% 7.65 68.85 63.63 50.96 49.38 30.79 52.72 ShortGPT 25% 7.15 69.67 65.19 51.26 52.46 33.89 54.49 LaCo 25% 7.38 68.53 63.76 50.39 51.28 33.45 53.48 Shortened LLaMA 25% 7.74 66.88 62.19 51.45 51.47 32.66 52.93 TRSP -$\ell_{2}$ 25% 6.53 71.35 67.62 55.84 52.26 35.75 56.56 TRSP -$\ell_{1}$ 25% 6.58 71.42 67.13 56.05 52.18 35.79 56.51 OPT-2.7B Dense 0% 12.46 74.81 61.01 60.58 54.42 31.14 56.39 SLEB 25% 15.71 65.19 56.26 44.54 46.28 25.14 47.48 ShortGPT 25% 14.96 67.37 57.65 46.82 49.43 26.54 49.56 LaCo 25% 15.38 66.54 59.45 43.68 48.74 24.96 48.67 Shortened LLaMA 25% 15.89 63.14 57.36 43.57 47.62 25.88 47.51 TRSP -$\ell_{2}$ 25% 13.18 70.54 60.27 46.35 51.59 27.36 51.22 TRSP -$\ell_{1}$ 25% 13.12 70.65 60.13 46.24 51.76 27.55 51.27 LLaMA2-7B Dense 0% 5.47 79.11 69.06 75.99 74.58 46.25 69.00 SLEB 25% 9.63 65.22 63.38 55.51 56.39 33.46 54.79 ShortGPT 25% 8.89 66.75 66.26 57.14 58.93 36.42 57.10 LaCo 25% 9.14 69.45 65.31 52.67 55.73 34.89 55.61 Shortened LLaMA 25% 9.47 65.58 64.72 58.36 54.19 32.96 55.16 TRSP -$\ell_{2}$ 25% 7.08 72.48 67.52 60.45 62.69 39.73 60.57 TRSP -$\ell_{1}$ 25% 7.17 72.08 67.93 60.42 62.38 39.89 60.54 LLaMA3-8B Dense 0% 5.76 85.56 77.94 79.27 78.84 56.49 75.62 SLEB 25% 10.38 72.74 64.12 67.74 65.84 45.16 63.12 ShortGPT 25% 9.26 75.38 69.25 70.12 68.54 47.57 66.17 LaCo 25% 10.14 74.69 67.52 66.36 69.43 46.31 64.86 Shortened LLaMA 25% 9.84 73.62 68.73 68.55 66.16 43.39 64.09 TRSP -$\ell_{2}$ 25% 7.84 77.25 71.63 72.26 71.49 49.58 68.44 TRSP -$\ell_{1}$ 25% 7.68 77.36 71.33 72.82 70.75 49.66 68.38 OPT-13B Dense 0% 10.12 76.82 64.80 69.81 61.87 35.67 61.79 SLEB 25% 11.96 72.83 64.06 63.32 59.98 34.65 58.97 ShortGPT 25% 11.38 73.59 64.52 65.68 60.41 34.99 59.84 LaCo 25% 11.79 73.96 63.24 62.17 61.46 33.28 58.82 Shortened LLaMA 25% 11.62 71.26 63.57 66.29 58.47 33.83 58.68 TRSP -$\ell_{2}$ 25% 10.45 74.15 64.47 68.82 61.55 35.23 60.84 TRSP -$\ell_{1}$ 25% 10.32 74.58 64.37 68.45 61.29 34.87 60.71 LLaMA2-13B Dense 0% 4.88 80.47 72.22 79.39 77.48 49.23 71.76 SLEB 25% 7.08 68.31 66.86 57.12 62.19 38.45 58.59 ShortGPT 25% 6.79 73.26 68.37 62.25 67.54 43.38 62.96 LaCo 25% 7.14 72.35 65.78 59.43 65.36 42.73 61.13 Shortened LLaMA 25% 6.92 69.45 66.38 59.86 65.25 39.12 60.01 TRSP -$\ell_{2}$ 25% 5.82 74.56 69.34 64.79 71.25 45.63 65.11 TRSP -$\ell_{1}$ 25% 5.89 75.06 69.18 64.96 71.43 45.26 65.18
@@
As shown in Table 1, TRSP achieves the lowest perplexity and the highest average accuracy across all models, demonstrating its superior performance on both generation and zero-shot tasks. Notably, on OPT-13B, TRSP only drop 1% in average accuracy compared to the dense model. On LLaMA2-7B, its perplexity is 20% lower than the second-best method, ShortGPT. This further demonstrates the effectiveness of TRSP. The minimal performance difference between $\ell_{2}$-norm and $\ell_{1}$-norm suggests that TRSP is not sensitive to the choice of regularization norm. In the following sections, we refer to using the $\ell_{2}$-norm in the stage two regularization.
@@ 4.3 Acceleration
LLMs language processing involves two key phases with different bottlenecks: compute-bound prompt processing and memory-bound token generation. We measured the speedup for each stage individually. Table 2 presents the latency and throughput results for OPT-13B and LLaMA2-13B running on a single 80GB NVIDIA A100 GPU. Following prior method  Song et al. (2024), the token generation was tested by producing 128-token sentences with a batch size of 64, and prompt processing latency was assessed with a 2048-token input.
Table 2: Throughput (tokens/s) and latency (ms) on OPT-13B and LLaMA2-13B. ‘PPL’ is the perplexity on Wikitext2. ‘PR’ is pruning ratio. ‘TI’ is throughput increase.
Model Method PR PPL($\downarrow$) Avg_Acc(%) Tokens/s TI Latency($\downarrow$) Speedup OPT-13B Dense 0% 10.12 61.79 1029 1.00× 386.5 1.00× TRSP 25% 10.45 60.84 1348 1.31× 286.3 1.35× TRSP 50% 15.38 50.83 1801 1.75× 208.9 1.85× LLaMA2-13B Dense 0% 4.88 71.76 1066 1.00× 396.9 1.00× TRSP 25% 5.82 65.11 1386 1.30× 298.4 1.33× TRSP 50% 11.28 57.57 1823 1.71× 216.9 1.83×
Pruning OPT-13B by 50% with TRSP yields a 75% increase in throughput and a 46% reduction in latency compared to the dense model. For LLaMA2-13B, it delivers a 71% improvement in throughput and a 45% decrease in latency. These results underscore the end-to-end acceleration achieved by TRSP.
@@ 4.4 Robustness to Different Pruning Ratios
Using the same settings as Section 4.2, we vary the pruning ratio from 20% to 60%. As shown in Figure 3, TRSP consistently achieves lower perplexity than other methods. At 20% pruning, it matches the dense model, and even at 60%, where SLEB fails, it maintains low perplexity. This demonstrates TRSP’s robustness and effectiveness in structured pruning for model acceleration. Additional results are provided in Appendix D.
4.5 Dependency on Datasets
Since TRSP relies on data-driven regularization, we investigate its dataset dependency. Keeping all other settings consistent with Section 4.2, we only change the source of the 128 samples: they are drawn respectively from Alpaca, WikiText-2, PTB, and C4. We evaluated perplexity of five methods on WikiText-2. As shown in Figure 4, TRSP consistently outperforms the other methods across datasets, demonstrating its robustness.
@@ 4.5 Dependency on Datasets
4.6 Minimal Cost
We only vary the retraining data size from 1,000 to 8,000 and use the same settings as in Section 4.2. We evaluated the perplexity of baselines on LLaMA2-7B. TRSP is retraining-free. From Figure 5, we observe that TRSP outperforms the other methods with 4,000 retraining data, significantly reducing the cost. We speculate that TRSP iteratively learns layer weights and then applies regularization, allowing it to identify layers to prune more accurately, and then transfer the knowledge from those layers to the remaining layers of the model through regularization. This process reduces knowledge loss, thus preserving model performance and lowering the retraining cost.
Figure 5: The perplexity on WikiText-2 using TRSP and baselines under different amounts of retraining data.
@@ 4.7 Ablation Study
Table 3: Ablation results on LLaMA2-7B and LLaMA2-13B. ‘w/o W’ denotes learning layer weights in one-shot, ‘w/o R’ means no regularization. The pruning ratio is 25%.
Model Setting PPL($\downarrow$) $\Delta$ AVG_ACC $\Delta$($\downarrow$) LLaMA2-7B TRSP 7.08 0.00 60.57 0.00 w/o W 9.26 +2.18 56.19 -4.38 w/o R 10.15 +3.07 54.36 -6.21 LLaMA2-13B TRSP 5.82 0.00 65.11 0.00 w/o W 8.35 +2.53 59.36 -5.75 w/o R 9.47 +3.65 56.25 -8.86
Effect of Learning Layer Weights Iteratively TRSP learns layer weights in a greedy and iterative manner. As shown in Table 3 (Row 3 and Row 6), replacing iterative layer weight learning with a one-shot approach leads to increased model perplexity and decreased accuracy, highlighting the importance of learning layer weights iteratively, which will be discussed in detail in Section 4.8.
@@
With all other settings the same as Section 4.2, we evaluate the model’s performance by varying $\lambda_{1}\in[10^{-5},10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}]$ and $\lambda_{2}\in[10^{-5},10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}]$. We perform a grid search over $\lambda_{1}$ and $\lambda_{2}$, with details provided in Appendix C. The optimal combination yielding the lowest perplexity on LLaMA2-7B is $\lambda_{1}=5\times 10^{-3}$ and $\lambda_{2}=10^{-3}$.
@@ 4.8 Choice of Learning Layer Weights
In this section, we explore the effectiveness of (1) iteratively learning layer weights using a greedy strategy compared to (2) acquiring all layer weights at one-shot. We set the pruning ratio to 25% and conduct experiments on LLaMA2-7B (32 layers), using the same settings as in Section 4.2.
@@
4.9 Impact of Stage Two Regularization
We keep other settings consistent with Section 4.2. After learning the layer weights iteratively, we consider two scenarios: (1) without regularization and (2) with regularization, then prune. The perplexity and average accuracy under (1) and (2) on LLaMA2-7B are in Table 5. The model exhibits lower perplexity and higher average accuracy with regularization. Since perplexity is the exponential form of cross-entropy loss, lower perplexity corresponds to a lower cross-entropy loss. A smaller loss function indicates better model performance indicating that the regularization process mitigates the impact of pruning on the overall model.
According to previous work Liu et al. (2023), the output and input of each layer in LLMs exhibit high similarity. This inherent similarity enables us to apply regularization on the difference between the input and output of certain layers using only a small portion of data. We compute the cosine similarity according to Equation 9 on the LLaMA2-7B model between the input representations $\mathbf{X_{\text{in}}^{i}}\in\mathbb{R}^{b\times n\times d}$ and the output representations $\mathbf{X_{\text{out}}^{i}}\in\mathbb{R}^{b\times n\times d}$, where $b$ denotes the batch size, $n$ the number of tokens, and $d$ is the hidden size of the model. The computation is performed on layers [3, 6, 9, 11, 13, 20, 23, 29], which are selected using the iterative method described in Section 4.8, both before and after regularization.
As shown in Figure 7, the input-output similarity of these layers is already high before regularization and increases even further after regularization. The increased similarity in the layers with regularization indicates that the input undergoes less change after passing through these layers, suggesting that less knowledge is retained in them. In the Appendix E, we illustrate the changes in similarity between the input and output of the layers that were not regularized, before and after regularization. The results in Figure 8 of Appendix E show that the similarity in these layers decreases, meaning that, after regularization, the input undergoes greater changes when passing through these layers than before. This demonstrates that the regularization process weakens the influence of the regularized layers and enhances the influence of the unregularized parts, suggesting that regularization process may facilitate the transfer of knowledge from the regularized layers to the rest of the model.
Table 5: The performance differences of (1) and (2).